FIRST HIGH-RESOLUTION SPECTRA AND ROTATIONAL CONSTANTS OF BrNO_{2} (NITRYL BRO-MIDE)

\underline{J. ORPHAL}, Institute of Environmental Physics, University of Bremen, P. O. Box 330440, 28334 Bremen, Germany; B. REDLICH, H. GROTHE, and H. WILLNER, Institute of Physical Chemistry, University of Hannover, Callinstr. 9, 30167 Hannover, Germany; A. FRENZEL and C. ZETZSCH, Fraunhofer-Institute of Toxicology and Aerosol Research, Fuchsstr. 1, 30625 Hannover, Germany.

Nitryl bromide (BrNO_{2}) formed by heterogeneous reactions on polar stratospheric clouds and on sea-salt particles in the marine troposphere. However, no rotationally resolved spectra have been reported up to now, and nothing is known about rotational constants, structure, and force field of BrNO_{2}. In fact, the only observations of BrNO_{2} in the gas phase were possible using flow experiments with gaseous N_{2}O_{5} flowing over NaBr particles. In this study, a different approach was used: By reaction of highly diluted gaseous ClNO_{2} with an aqueous Br^{−} solution, BrNO_{2} was formed and trapped at low temperature. The crude product was purified by trap-to-trap condensations and fractional sublimation in vacuo. High-resolution infrared spectra of the ν_{4} band of BrNO_{2} around 1660 cm^{−1} were measured with resolutions of up to 0.002 cm^{−1}, using the Bruker IFS-120HR FTS at University of Hannover. For the first time, rotational constants for the 79BrNO_{2} and ^{81}BrNO_{2} isotopomers were determined for the ν_{4} and ground vibrational states. The constants were used for determination of a first substitution structure and, together with infrared band centers, the harmonic force field of BrNO_{2}.